[figure: 9.7 cm]

KONDO LATTICE STATES WITH COMPENSATED MOMENTS

AND SPIN DEPENDENT HEAVY MASSES

J. Spałek

Marian Smoluchowski Institute of Physics, Jagiellonian University, ulica Reymonta 4, 30-059 Kraków, Poland

I review a recent work [1,2] on the magnetic Kondo-lattice metallic and insulating states, for which the almost compensated moments have been obtained self-consistently for the first time. This heavy-electron phase can border with a ferromagnetic phase with an equally weak resultant moment. Both of these phases have (an almost) integer valency of f electrons, but only the first of them has a very large density of states at the Fermi level. The Kondo insulating gap is strongly reduced by the thermal fluctuations and disappears in a noncritical manner. We also show that a consistent mean-field theory of correlated electrons is flawed unless the heavy masses of quasiparticles are spin-dependent in the polarized state.

[1]. R. Doradziński and J. Spałek, Phys. Rev. B65, R14239 (1997); B58, 3293 (1998); and unpublished. [2]. J. Spałek, unpublished.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : PLENARY LECTURE

Corresponding author : Jozef Spałek

Address for correspondence : Marian Smoluchowski Institute of Physics, Jagiellonian University, ulica Reymonta 4, 30-059 Kraków, Poland

Email address : ufspalek@jetta.if.uj.edu.pl